OpenStudy (anonymous):

Kevin plans to put concrete on a rectangular portion of his driveway. The portion is 10 feet long and 3 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Kevin needs is $72.59. Which of the following is closest to the width of the portion of the driveway on which Kevin plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet] 3 feet 6 feet 8 feet 10 feet

OpenStudy (anonymous):

@Melissa_Something

OpenStudy (melissa_something):

I'm trippin, I'm tryna set up a proportion but this making me mad hahaha

OpenStudy (anonymous):

lol is there anyway you can help me with this one

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@surry99 are you going to help me?

OpenStudy (surry99):

Convert 10 feet to yards = x Convert 3 inches to yards = y width in yard = z So xyz($98/yd^3) = $72.59 Solve for z in yards then convert it to feet

OpenStudy (anonymous):

8 feet

OpenStudy (anonymous):

you there

OpenStudy (surry99):

yep

OpenStudy (anonymous):

that is the correct answer

OpenStudy (surry99):

i havent done the math

OpenStudy (anonymous):

oh

OpenStudy (surry99):

plug the numbers in and check to see if it works

OpenStudy (surry99):

Does it work when you check?

OpenStudy (anonymous):

idk

OpenStudy (surry99):

hang on...

OpenStudy (anonymous):

please hurry can you also help me with another one after this

OpenStudy (ccswims):

I got 8 feet

OpenStudy (anonymous):

can you help me with another one

OpenStudy (ccswims):

I can try

OpenStudy (anonymous):

ok

OpenStudy (ccswims):

Do you take FLVS cause I remember a question like this on an exam

OpenStudy (anonymous):

The length of the base of a triangle is 5.3 cm and its height is 6.7 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2. (Round your answer to the appropriate significant digits) Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

It is for a practice test

OpenStudy (surry99):

.5 (5.3 cm) x 6.7 cm =

OpenStudy (ccswims):

1/2 (5.3 * 6.7)

OpenStudy (anonymous):

35.51 is the answer

OpenStudy (ccswims):

that gives you 1/2 * 35.51

OpenStudy (ccswims):

no now you must divide by 1/2 again

OpenStudy (anonymous):

17.755

OpenStudy (ccswims):

Round it and your done :)

OpenStudy (anonymous):

but i have to round to significant figures which 17.8 and i got that wrong on the pretest

OpenStudy (ccswims):

Does it say where to round it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it says: round your answer to the appropriate significant figures

OpenStudy (anonymous):

are you still here?

OpenStudy (ccswims):

I would just type 31.5 then

OpenStudy (anonymous):

why

OpenStudy (anonymous):

that might be the wrong answer

OpenStudy (ccswims):

Honestly I truly believe it is 17.8, but when it says "appropriate figures" it might mean 17.755 like 17.76

OpenStudy (anonymous):

that might be it

OpenStudy (anonymous):

i will put it in may I have help with a few more

OpenStudy (anonymous):

@ccswims

OpenStudy (ccswims):

Sure

OpenStudy (anonymous):

An equation is shown below: -2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation? The equation can become -2(4x - 1) = -2 by applying the distributive property. The equation can become -2(4x - 1) = 12 by applying the addition property of equality. The equation can become -2(4x - 1) = 12 by applying the commutative property of addition The equation can become -2(4x - 1) = -2 by applying the subtraction property of equality.

OpenStudy (anonymous):

I am timed on this test for some reason may we go a little quicker

OpenStudy (ccswims):

B

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

@ccswims

OpenStudy (ccswims):

You cancel out -7 to 7 and add 12 + 7

OpenStudy (ccswims):

this is the addition property of equality

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

may you help me with two more

OpenStudy (ccswims):

yea

OpenStudy (anonymous):

A bulldozer is rented at $800 per day plus a charge per hour of use. The bulldozer was used for 5 hours on a day and the total charge was $3000. Which equation can be used to find the charge x, in dollars, per hour of use?

OpenStudy (anonymous):

800 + 5x = 3000 800x + 5 = 3000 800x + 5x = 3000 800 + 5 = 3000x

OpenStudy (ccswims):

A because x is the rent per hour

OpenStudy (anonymous):

A car repair shop charged $80 per hour for labor and $210 for the parts of a car. If the total cost of repair for the car was $490, how many hours did it take to repair the car? 0.59 2.71 3.50 8.75

OpenStudy (ccswims):

C

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i am submitting it lets hope I get a good grade

OpenStudy (anonymous):

@ccswims I got one wrong

OpenStudy (ccswims):

Which one?

OpenStudy (anonymous):

it was the one about rounding the one that we could not understand

OpenStudy (ccswims):

I'm sorry, honestly I believe it was 17.8 because they said round

OpenStudy (anonymous):

that is okay though I am still happy I got a 95

OpenStudy (ccswims):

But you said that was wrong

OpenStudy (anonymous):

i know

OpenStudy (ccswims):

Good :) I'm really sorry about that one

OpenStudy (anonymous):

I am going to email my teacher and ask him the correct answer

OpenStudy (ccswims):

Yea, cause that one was confusing because it also didn't tell you to round to the nearest tenth or hundredth

OpenStudy (anonymous):

exactly, but thanks for all the help you are awesome i am going to fan you